# Estimation of partially known Gaussian graphical models with score-based structural priors

### 25 Jan 2024 v1

@@ 1 Introduction
All techniques proposed so far for GGM estimation impose limitations in the prior knowledge that can be incorporated. Importantly, the imposed priors are too simple (e.g., sparsity) or restrictive (e.g., G-Wishart), and these are imposed on $\boldsymbol{\Theta}$ while structural priors would be more naturally defined on the underlying adjacency matrix ${\mathbf{A}}$.
Addressing these limitations. This paper proposes a new approach that allows the introduction of arbitrary prior information directly on ${\mathbf{A}}$, based on a dataset of adjacency matrices ${\mathcal{A}}$ (of potentially varying sizes) whose distribution we use as a prior $p({\mathbf{A}})$. This is particularly useful for real-world applications where we often have datasets of graphs instead of a closed-form prior distribution $p({\mathbf{A}})$. For instance, when learning brain networks, leveraging graphs from other patients is feasible (and valuable) because they often share similar structures. This idea also applies to other areas like molecular datasets or social networks. Additionally, generating samples to create a synthetic ${\mathcal{A}}$ is easy for many random graphs, but the associated graph distribution cannot be expressed in closed form. Relevant examples include Barabási–Albert and small world graphs. In the particular case of the actual distribution $p({\mathbf{A}})$ being known, one could sample from it and generate the dataset of graphs ${\mathcal{A}}$.
Additionally, we allow edge-to-edge constraints (in a WGL fashion) so that edges that are known to either exist or be absent lead to values of $1$ or 00 in ${\mathbf{A}}$, respectively. This is useful in cases where some pairs of variables are known to be conditionally independent, but we want to estimate the rest of the graph. Examples include gene expression data (Li and Jackson,, 2015), neurotoxicology tests (Grzebyk et al.,, 2004), social networks (Wu et al.,, 2019), or brain networks (Simpson and Laurienti,, 2015).
@@ 2 Problem formulation
Throughout this work, we refer to the known and unknown fractions of the adjacency matrix as ${\mathbf{A}}^{\mathcal{O}}$ and ${\mathbf{A}}^{\mathcal{U}}$, respectively. The condition that $i<j$ implies that we do not take into account the diagonal (since the graph has no self-loops), and we just consider the upper-triangular part of ${\mathbf{A}}$ (since it is symmetric).
@@ 3 Langevin for support estimation
This section explains how to use annealed Langevin dynamics to solve Problem 1. In Section 3.1, we propose a distribution: i) that approximates the actual posterior (8) and ii) from which samples can be drawn by leveraging Langevin dynamics. Based on this approximate posterior, we define an estimator of ${\mathbf{A}}_{0}$ and show that it is consistent. Sections 3.2 and 3.3 explain how Langevin dynamics and its annealed counterpart work and why they provide a way to incorporate prior information in the estimation of the binary matrix ${\mathbf{A}}_{0}$ via the so-called score function. Then, in Section 3.4, we study how to use the dataset ${\mathcal{A}}$ as prior knowledge by training a graph neural network (GNN) whose output is directly plugged into the Langevin dynamics. Section 3.5 describes our final algorithm, which combines and summarizes the results of this section.
@@ 3.1 Proposed estimator
which corresponds to the positive definite matrix that maximizes the likelihood while respecting the zero pattern known to exist. The optimization problem in (3.1) can be efficiently solved by using the WGL algorithm described in Section 1. The constraint is equivalent to setting a penalty $\Lambda_{ij}$ to an arbitrarily large constant for those entries where ${\mathbf{A}}$ is known to be 00, and setting $\Lambda_{ij}=0$ otherwise.
@@
where we recall that $\circ$ is the entry-wise product. Since the entries of $({\mathbf{A}}+{\mathbf{I}})$ are binary, the entry-wise multiplication can be understood as a mask that sets to zero the entries of the precision that are not associated with an edge. Let us suppose now that we can sample from (10) and let $\{{\mathbf{A}}^{(m)}\}_{m=1}^{M}$ denote the set of $M$ generated independent samples. Then, the set $\{{\mathbf{A}}^{(m)}\}_{m=1}^{M}$ can be used to characterize the posterior. We focus on the posterior sample mean estimator presented in (6)–(7). Then, the estimator for ${\mathbf{A}}_{0}$ that we propose boils down to
@@
See Section F.1 in the Supplementary Material (SM). ∎
@@ 3.3 Annealed Langevin dynamics
The annealed version of the dynamics was initially introduced to allow the algorithm to converge faster and perform better (Song and Ermon,, 2019). However, in our case, it also offers the advantage of rendering the problem differentiable. Consequently, the annealing enables the computation of the score functions and the use of Langevin dynamics to sample from an originally discrete distribution. If the noise levels $\{\sigma_{l}\}_{l=1}^{L}$ and the step size $\epsilon$ are chosen adequately (Song and Ermon,, 2019), after a sufficiently large number of iterations, the sample ${\tilde{\mathbf{a}}}_{t}$ is arbitrarily close to an actual sample from the discrete distribution $p({\mathbf{a}}\mid{\mathbf{X}})$. The noisy sample ${\tilde{\mathbf{a}}}_{t}$ must be projected onto the set $\{0,1\}$ if an actual sample is needed.
@@ 3.4 Learned annealed scores
Let ${\mathbf{g}}_{\boldsymbol{\xi}}({\tilde{\mathbf{a}}},\sigma)$ be the output of the GNN we wish to train, with $\boldsymbol{\xi}$ being its trainable parameters. Ideally, the output for a given ${\tilde{\mathbf{a}}}$ (with the associated noise level $\sigma_{l}$ of the current iteration) should be as close as possible to the actual score $\nabla\log p({\tilde{\mathbf{a}}})$. The loss function to learn $\boldsymbol{\xi}$ should be designed to minimize the mean squared error across all noise levels jointly. To achieve this, we define the distance
@@ 4 Numerical results
When $|{\mathcal{U}}|$ is small, the prior probability mass is concentrated among fewer possible graphs. Intuitively, in this case, Langevin generally samples either the same graph or similar ones throughout the different $M$ samples. Thus, the sample mean yields a satisfactory estimate. When $|{\mathcal{U}}|$ is large, the probability mass is spread across many adjacency matrices in the high-dimensional space of $p({\tilde{\mathbf{A}}})$. This leads to Langevin converging to diverse graphs each time we sample, reducing the usefulness of the sample mean as an estimator. An additional experiment illustrating the performance dependence on $|{\mathcal{U}}|$ is presented in Section H.2 of the SM.
@@
The F1 score achieved by LPr is relatively poor due to the large randomness in the generation of the underlying graphs. Namely, it is always worse than the one obtained using WGL. On the contrary, when the underlying graph presents more structure (for instance, the grid graphs in Figure 0(a)), LPr was shown to outperform WGL for some values of $k/|{\mathcal{U}}|$. We can conclude that some priors offer more predictive power than others: the more substantial the structure of the graphs, the more useful $p({\mathbf{A}})$ becomes.
@@
Ego-nets, like grids, present a strong structure, rendering the prior highly predictive. Even though half of the graph is unknown, the F1 score of LPr is the highest among all the experiments (cf. Figures 1 and 1(a)).
@@
Overall, our numerical experiments show that i) our approach leads to better graph estimation results than the classical alternatives considered and ii) the benefits of our approach are more significant when the number of observations is small, and the graph presents marked structural features.
@@ Acknowledgments
F PROOFS OF LEMMAS
F.1 Proof of Lemma 1
@@
$\hat{\boldsymbol{\Theta}}=h({\mathbf{S}}),$ (SM31)
@@
First, we compute $h\left(\boldsymbol{\Theta}^{-1}_{0}\right)$. To this end, we first check what matrix maximizes $f$ without considering the constraint in (F.1). Since $f$ is continuous and strictly concave in ${\mathcal{S}}_{+}^{n}$ (Ravikumar et al.,, 2011), if follows that the maximizer is unique. Taking the gradient of $f$ with respect to $\boldsymbol{\Theta}$ yields
$\frac{\partial f(\boldsymbol{\Theta};{\mathbf{V}})}{\partial\boldsymbol{\Theta}}=2\boldsymbol{\Theta}^{-1}-\boldsymbol{\Theta}^{-1}\circ{\mathbf{I}}-2{\mathbf{V}}-{\mathbf{V}}\circ{\mathbf{I}}=\boldsymbol{0}\iff\boldsymbol{\Theta}={\mathbf{V}}^{-1}.$ (SM32)
Hence, $f(\boldsymbol{\Theta};\boldsymbol{\Theta}_{0}^{-1})$ is maximized when $\boldsymbol{\Theta}=\boldsymbol{\Theta}_{0}$. Furthermore, $\boldsymbol{\Theta}_{0}$ satisfies the constraint in (F.1). As a result, it holds that
$h\left(\boldsymbol{\Theta}^{-1}_{0}\right)=\boldsymbol{\Theta}_{0}.$ (SM33)
Notice that the mapping $h$ is continuous. This can be proven by means of the maximum theorem (Sundaram,, 1996). Let ${\mathcal{C}}=\{{\mathbf{V}}\in{\mathbb{R}}^{n\times n}\mid V_{ij}=0\,\,\,\,\forall(i,j):A^{\mathcal{O}}_{ij}=0\}$ be the set of constrained matrices we are interested in. Any linear combination of matrices in ${\mathcal{C}}$ is still in ${\mathcal{C}}$, and thus ${\mathcal{C}}$ is a convex set. Therefore, since ${\mathcal{S}}_{+}^{n}$ is convex as well, the intersection ${\mathcal{I}}={\mathcal{C}}\cap{\mathcal{S}}_{+}^{n}$ (which is the set over which $f$ is maximized in (F.1) to compute $h$) is convex too.
@@
${\mathbf{S}}=\frac{1}{k}\sum_{i=1}^{k}{\mathbf{x}}_{i}{\mathbf{x}}_{i}^{\top}\xrightarrow{k\to\infty}{\mathbb{E}}\left[{\mathbf{x}}{\mathbf{x}}^{\top}\right]=\boldsymbol{\Theta}_{0}^{-1}.$ (SM34)
Consequently, by the continuous mapping theorem (which can be applied because $h$ is continuous), considering (SM34), (SM31) and (SM33), we conclude that
$\hat{\boldsymbol{\Theta}}=h({\mathbf{S}})\xrightarrow{k\to\infty}h\left(\boldsymbol{\Theta}_{0}^{-1}\right)=\boldsymbol{\Theta}_{0}.$ (SM35)
F.2 Proof of Lemma 2
In order to study how ${\hat{p}}({\mathbf{A}}\mid{\mathbf{X}})$ behaves as $k\to\infty$, we first analyze the likelihood function ${\mathcal{L}}_{\mathbf{X}}\left(\boldsymbol{\Theta}\right)$. Using Bayes’ theorem, we get
$p(\boldsymbol{\Theta}\mid{\mathbf{X}})\propto{\mathcal{L}}_{\mathbf{X}}\left(\boldsymbol{\Theta}\right)p(\boldsymbol{\Theta}).$ (SM36)
By the Bernstein–von Mises theorem, we know that $p(\boldsymbol{\Theta}\mid{\mathbf{X}})\xrightarrow{k\to\infty}\delta(\boldsymbol{\Theta}-\boldsymbol{\Theta}_{0})$, where $\delta(\cdot)$ is the Dirac delta. Combining this with (SM36) it follows that
${\mathcal{L}}_{\mathbf{X}}\left(\boldsymbol{\Theta}\right)\xrightarrow{k\to\infty}0\quad\forall\boldsymbol{\Theta}\neq\boldsymbol{\Theta}_{0}.$ (SM37)
The result in (SM37) holds for any prior $p(\boldsymbol{\Theta})$, since the prior can be considered a constant with respect to $k$.
Given (SM37) and the consistency of $\hat{\boldsymbol{\Theta}}$ in Lemma 1, we conclude that [cf. (10) in the main paper]
${\hat{p}}({\mathbf{A}}\mid{\mathbf{X}})\xrightarrow{k\to\infty}\frac{p({\mathbf{A}})}{C}\delta\left(\boldsymbol{\Theta}_{0}\circ({\mathbf{A}}+{\mathbf{I}})-\boldsymbol{\Theta}_{0}\right),$ (SM38)

### 23 Feb 2024 v3

@@ 1 INTRODUCTION
All techniques proposed so far for GGM estimation impose limitations in the prior knowledge that can be incorporated. Importantly, the imposed priors are too simple (e.g., sparsity) or restrictive (e.g., G-Wishart), and these are imposed on $\boldsymbol{\Theta}$ while the natural approach would be to impose the structural priors on the underlying adjacency matrix ${\mathbf{A}}$.
Addressing these limitations. This paper proposes a new approach that allows the introduction of arbitrary prior information directly on ${\mathbf{A}}$, based on a dataset of adjacency matrices ${\mathcal{A}}$ (of potentially varying sizes) whose distribution we use as a prior $p({\mathbf{A}})$. This is particularly useful for real-world applications where we often have datasets of graphs instead of a closed-form prior distribution $p({\mathbf{A}})$. For instance, when learning brain networks, leveraging graphs from other patients is feasible (and valuable) because they often share similar structures. This idea also applies to other areas like molecular datasets or social networks. Additionally, many random graphs do not have a closed-form distribution, but generating samples to create a synthetic ${\mathcal{A}}$ is relatively easy.
Furthermore, we allow edge-to-edge constraints (in a WGL fashion) so that edges that are known to either exist or be absent lead to values of $1$ or 00 in ${\mathbf{A}}$, respectively. This is useful in cases where some pairs of variables are known to be conditionally independent, but we want to estimate the rest of the graph. Examples include gene expression data (Li and Jackson,, 2015), neurotoxicology tests (Grzebyk et al.,, 2004), social networks (Wu et al.,, 2019), or brain networks (Simpson and Laurienti,, 2015).
@@ 2 PROBLEM FORMULATION
Throughout this work, we refer to the known and unknown fractions of the adjacency matrix as ${\mathbf{A}}^{\mathcal{O}}$ and ${\mathbf{A}}^{\mathcal{U}}$, respectively. The condition that $i<j$ implies that we do not take into account the diagonal (since the graph has no self-loops), and we just consider the upper-triangular part of ${\mathbf{A}}$ (since it is symmetric). In case the set of edges is completely unknown, we can estimate some of its entries with high confidence and consider them known, as we show in Section 4.
@@ 3 LANGEVIN FOR SUPPORT ESTIMATION
This section explains how to use annealed Langevin dynamics to solve Problem 1. In Section 3.1, we propose a distribution: i) that approximates the actual posterior (8) and ii) from which samples can be drawn by leveraging Langevin dynamics. Based on this approximate posterior, we define an estimator of ${\mathbf{A}}_{0}$ and show that it is consistent. Sections 3.2 and 3.3 explain how annealed Langevin dynamics works and why it provides a way to incorporate prior information in the estimation ${\mathbf{A}}_{0}$ via the so-called score function. Then, in Section 3.4, we study how to use the dataset ${\mathcal{A}}$ as prior knowledge by training a graph neural network (GNN) whose output is directly plugged into the Langevin dynamics. Section 3.5 describes our final algorithm, which combines and summarizes the results of this section. An illustration of the overall procedure is shown in Figure 1.
Figure 1: Illustration of our final algorithm [cf. Algorithm 1]. The grey entries in ${\mathbf{A}}^{{\mathcal{O}}}_{0}$ are what we aim to estimate (i.e., ${\mathbf{A}}^{{\mathcal{U}}}$). If no entries of ${\mathbf{A}}_{0}$ are known, some can be estimated by bootstrapping ${\mathbf{X}}$, as shown in Section 4. By combining the GGM observations ${\mathbf{X}}$ and the partially observed graph ${\mathbf{A}}^{{\mathcal{O}}}_{0}$, we compute the constrained ML estimator $\hat{\boldsymbol{\Theta}}$ by solving (3.1). This encodes information about the likelihood of ${\mathbf{A}}$ given ${\mathbf{X}}$. To encode information about the prior $p({\mathbf{A}})$, we process the dataset ${\mathcal{A}}$ with a GNN (Section 3.4). Then, we can draw $M$ samples from the (approximate) posterior using a Langevin sampler and build any estimator with them, such as ${\hat{\mathbf{A}}}$ in (11) that approximates the posterior mean.
@@ 3.1 Proposed estimator
which corresponds to the positive definite matrix that maximizes the likelihood while respecting the zero pattern known to exist. The optimization problem in (3.1) can be efficiently solved by using the WGL algorithm mentioned in Section 1. The constraint is equivalent to setting a penalty $\Lambda_{ij}$ to an arbitrarily large constant for those entries where ${\mathbf{A}}$ is known to be 00, and setting $\Lambda_{ij}=0$ otherwise.
@@
where we recall that $\circ$ is the entry-wise product. Since the entries of $({\mathbf{A}}+{\mathbf{I}})$ are binary, the entry-wise multiplication can be understood as a mask that sets to zero the entries of the precision that are not associated with an edge. Let us suppose now that we can sample from (10) and let $\left\{{\mathbf{A}}^{(m)}\right\}_{m=1}^{M}$ denote the set of $M$ generated independent samples. Then, the set $\left\{{\mathbf{A}}^{(m)}\right\}_{m=1}^{M}$ can be used to characterize the posterior. We focus on the posterior sample mean estimator presented in (6)–(7). Then, the estimator for ${\mathbf{A}}_{0}$ that we propose boils down to
@@
See Section A.1 in the Supplementary Material (SM). ∎
@@
In practical scenarios, infinite samples are never available, yet consistency is a desirable property for an estimator. Furthermore, even though Theorem 1 requires $M\to\infty$, our experiments (Section 4) reveal that our method outperforms classical methods for relatively small values of $M$.
@@ 3.3 Annealed Langevin dynamics
The annealed version of the dynamics was initially introduced to allow the algorithm to converge faster and perform better (Song and Ermon,, 2019). However, in our case, it also offers the advantage of rendering the problem differentiable. Consequently, the annealing enables the computation of the score functions and the use of Langevin dynamics to sample from an originally discrete distribution. If the noise levels $\{\sigma_{l}\}_{l=1}^{L}$ and the step size $\epsilon$ are chosen adequately (Song and Ermon,, 2019), after a sufficiently large number of iterations, the sample ${\tilde{\mathbf{a}}}_{t}$ is arbitrarily close to an actual sample from the discrete distribution $p({\mathbf{a}}\mid{\mathbf{X}})$. If an actual sample is needed, the noisy sample ${\tilde{\mathbf{a}}}_{t}$ must be projected onto the set $\{0,1\}$.
@@ 3.4 Learned annealed scores
Let ${\mathbf{g}}_{\boldsymbol{\xi}}({\tilde{\mathbf{a}}},\sigma)$ be the output of the GNN we wish to train, with $\boldsymbol{\xi}$ being its trainable parameters. Ideally, the output for a given ${\tilde{\mathbf{a}}}$ (with the associated noise level $\sigma_{l}$ of the current iteration) should be as close as possible to the actual score $\nabla\log p({\tilde{\mathbf{a}}})$. The loss function to learn $\boldsymbol{\xi}$ should be designed to jointly minimize the mean squared error across all noise levels. To achieve this, we define the distance
@@ 4 NUMERICAL RESULTS
When $|{\mathcal{U}}|$ is small, the prior probability mass is concentrated among fewer possible graphs. Intuitively, in this case, Langevin generally samples either the same graph or similar ones throughout the different $M$ samples. Thus, the sample mean yields a satisfactory estimate. When $|{\mathcal{U}}|$ is large, the probability mass is spread across many adjacency matrices in the high-dimensional space of $p({\tilde{\mathbf{A}}})$. This leads to Langevin converging to diverse graphs each time we sample, reducing the usefulness of the sample mean as an estimator. An additional experiment illustrating the performance dependence on $|{\mathcal{U}}|$ is presented in Section C.2 of the SM.
@@
The F1 score achieved by LPr is relatively poor due to the large randomness in the graphs. Namely, it is always worse than the one obtained using WGL. On the contrary, when the underlying graph presents more structure (for instance, the grid graphs in Figure 1(a)), LPr was shown to outperform WGL for some values of $k/|{\mathcal{U}}|$. We can conclude that some priors offer more predictive power than others: the more substantial the structure of the graphs, the more useful $p({\mathbf{A}})$ becomes.
@@
Ego-nets, like grids, present a strong structure, rendering the prior highly predictive. Even though half of the graph is unknown, the F1 score of LPr is the highest among all the experiments (cf. Figures 2 and 2(a)).
@@
Fully unknown ego-nets. We consider the same dataset as in the last experiment, but now all the entries in ${\mathbf{A}}$ (except for those in the diagonal, which are 00) are assumed unknown so that $|{\mathcal{O}}|=0$. As shown in Figure 2, the prior offers less predictive power as $|{\mathcal{U}}|$ increases (see Section C.2 for another experiment investigating this behavior). Thus, when the entire graph is to be estimated, we propose fixing some of the entries in ${\mathbf{A}}$ with a graphical version of the random lasso (Wang et al.,, 2011).
To that end, we first compute the GL solution $B$ times (we use $B=50$ in these experiments), where a different bootstrap sample ${\mathbf{X}}^{(b)}$ is used for each iteration to obtain $\hat{\boldsymbol{\Theta}}^{(b)}_{\mathrm{boot}}$. Then, the average
${\hat{\mathbf{A}}}_{\mathrm{boot}}=\frac{1}{B}\sum_{b=1}^{B}{\mathbb{I}\left\{\hat{\boldsymbol{\Theta}}^{(b)}_{\mathrm{boot}}\neq 0\right\}}$ (30)
can be interpreted as the probability of each $A_{ij}$ to be $1$. Thus, for some probability margin $p_{m}$, we assume known some entries $A_{ij}$ such that
$A^{{\mathcal{O}}}_{ij}=\begin{cases}0&\text{if }{\hat{A}}_{\mathrm{boot}_{ij}}<0.5-p_{m}\\ 1&\text{if }{\hat{A}}_{\mathrm{boot}_{ij}}>0.5+p_{m}\end{cases},$ (31)
leaving as unknown all entries $(i,j)$ such that $0.5-p_{m}\leq{\hat{A}}_{\mathrm{boot}_{ij}}\leq 0.5+p_{m}$. As $p_{m}$ increases, the confidence of the estimated fixed values is higher, and $|{\mathcal{O}}|$ becomes smaller.
The results for two different values of $p_{m}$ are shown in Figure 4, where we also compare with GL and its bootstrapped counterpart (BGL). The latter is computed as in (30) and then thresholded. We observe that a naive implementation of our method falls behind when $|{\mathcal{O}}|=0$, an expected behavior as analyzed in the experiments with grid graphs. However, by leveraging the bootstrapping procedure to fix some entries in ${\mathbf{A}}$ we outperform GL and BGL, indicating that the prior distribution significantly contributes to the prediction accuracy.
Overall, our numerical experiments show that i) our approach leads to better graph estimation results than the classical alternatives considered and ii) the benefits of our approach are more significant when the number of observations is small and the graph presents marked structural features.
Figure 4: F1 score comparison when estimating ego-nets with no known values in ${\mathbf{A}}$. The values in parentheses correspond to the $p_{m}$ used to fix values from ${\hat{\mathbf{A}}}_{\mathrm{boot}}$ prior to using our method LPost.
@@ Acknowledgments
Checklist
  1. 1.
    For all models and algorithms presented, check if you include:
    1. (a)
      A clear description of the mathematical setting, assumptions, algorithm, and/or model. Yes.
    2. (b)
      An analysis of the properties and complexity (time, space, sample size) of any algorithm. Yes. Additional analysis is provided in the SM. In particular, for a complexity analysis please refer to Section B.3.
    3. (c)
      (Optional) Anonymized source code, with specification of all dependencies, including external libraries. Yes.
  2. 2.
    For any theoretical claim, check if you include:
    1. (a)
      Statements of the full set of assumptions of all theoretical results. Yes.
    2. (b)
      Complete proofs of all theoretical results. Yes. Proofs to the auxiliary lemmas are provided in the SM (Section A).
    3. (c)
      Clear explanations of any assumptions. Yes.
  3. 3.
    For all figures and tables that present empirical results, check if you include:
    1. (a)
      The code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL). Yes. Please refer to the source code if needed.
    2. (b)
      All the training details (e.g., data splits, hyperparameters, how they were chosen). Yes. Please refer to Sections B.1 and B of the SM for more information.
    3. (c)
      A clear definition of the specific measure or statistics and error bars (e.g., with respect to the random seed after running experiments multiple times). Yes. Additional information can be found in Section B of the SM.
    4. (d)
      A description of the computing infrastructure used. (e.g., type of GPUs, internal cluster, or cloud provider). Yes. See Section B.3 in the SM.
  4. 4.
    If you are using existing assets (e.g., code, data, models) or curating/releasing new assets, check if you include:
    1. (a)
      Citations of the creator if your work uses existing assets. Yes.
    2. (b)
      The license information of the assets, if applicable. Yes. Publication of our source code is available in GitHub under an MIT License.
    3. (c)
      New assets either in the supplemental material or as a URL, if applicable. Yes.
    4. (d)
      Information about consent from data providers/curators. Not Applicable.
    5. (e)
      Discussion of sensible content if applicable, e.g., personally identifiable information or offensive content. Not Applicable.
  5. 5.
    If you used crowdsourcing or conducted research with human subjects, check if you include:
    1. (a)
      The full text of instructions given to participants and screenshots. Not Applicable
    2. (b)
      Descriptions of potential participant risks, with links to Institutional Review Board (IRB) approvals if applicable. Not Applicable
    3. (c)
      The estimated hourly wage paid to participants and the total amount spent on participant compensation. Not Applicable
Estimation of partially known Gaussian graphical models with score-based structural priors
@@
A PROOFS OF LEMMAS
A.1 Proof of Lemma 1
@@
$\hat{\boldsymbol{\Theta}}=h({\mathbf{S}}),$ (33)
@@
First, we compute $h\left(\boldsymbol{\Theta}^{-1}_{0}\right)$. To this end, we first check what matrix maximizes $f$ without considering the constraint in (A.1). The maximizer is unique since $f$ is continuous and strictly concave in ${\mathcal{S}}_{+}^{n}$ (Ravikumar et al.,, 2011). Taking the gradient of $f$ with respect to $\boldsymbol{\Theta}$ yields
$\frac{\partial f(\boldsymbol{\Theta};{\mathbf{V}})}{\partial\boldsymbol{\Theta}}=2\boldsymbol{\Theta}^{-1}-\boldsymbol{\Theta}^{-1}\circ{\mathbf{I}}-2{\mathbf{V}}-{\mathbf{V}}\circ{\mathbf{I}}=\boldsymbol{0}\iff\boldsymbol{\Theta}={\mathbf{V}}^{-1}.$ (34)
Hence, $f(\boldsymbol{\Theta};\boldsymbol{\Theta}_{0}^{-1})$ is maximized when $\boldsymbol{\Theta}=\boldsymbol{\Theta}_{0}$. Furthermore, $\boldsymbol{\Theta}_{0}$ satisfies the constraint in (A.1). As a result, it holds that
$h\left(\boldsymbol{\Theta}^{-1}_{0}\right)=\boldsymbol{\Theta}_{0}.$ (35)
Notice that the mapping $h$ is continuous. This can be proven through the maximum theorem (Sundaram,, 1996). Let ${\mathcal{C}}=\{{\mathbf{V}}\in{\mathbb{R}}^{n\times n}\mid V_{ij}=0\,\,\,\,\forall(i,j):A^{\mathcal{O}}_{ij}=0\}$ be the set of constrained matrices we are interested in. Any linear combination of matrices in ${\mathcal{C}}$ is still in ${\mathcal{C}}$; thus, ${\mathcal{C}}$ is a convex set. Therefore, since ${\mathcal{S}}_{+}^{n}$ is convex as well, the intersection ${\mathcal{I}}={\mathcal{C}}\cap{\mathcal{S}}_{+}^{n}$ (which is the set over which $f$ is maximized in (A.1) to compute $h$) is convex too.
@@
${\mathbf{S}}=\frac{1}{k}\sum_{i=1}^{k}{\mathbf{x}}_{i}{\mathbf{x}}_{i}^{\top}\xrightarrow{k\to\infty}{\mathbb{E}}\left[{\mathbf{x}}{\mathbf{x}}^{\top}\right]=\boldsymbol{\Theta}_{0}^{-1}.$ (36)
Consequently, by the continuous mapping theorem (which can be applied because $h$ is continuous), considering (36), (33) and (35), we conclude that
$\hat{\boldsymbol{\Theta}}=h({\mathbf{S}})\xrightarrow{k\to\infty}h\left(\boldsymbol{\Theta}_{0}^{-1}\right)=\boldsymbol{\Theta}_{0}.$ (37)
A.2 Proof of Lemma 2
To study how ${\hat{p}}({\mathbf{A}}\mid{\mathbf{X}})$ behaves as $k\to\infty$, we first analyze the likelihood function ${\mathcal{L}}_{\mathbf{X}}\left(\boldsymbol{\Theta}\right)$. Using Bayes’ theorem, we get
$p(\boldsymbol{\Theta}\mid{\mathbf{X}})\propto{\mathcal{L}}_{\mathbf{X}}\left(\boldsymbol{\Theta}\right)p(\boldsymbol{\Theta}).$ (38)
By the Bernstein–von Mises theorem, we know that $p(\boldsymbol{\Theta}\mid{\mathbf{X}})\xrightarrow{k\to\infty}\delta(\boldsymbol{\Theta}-\boldsymbol{\Theta}_{0})$, where $\delta(\cdot)$ is the Dirac delta. Combining this with (38) it follows that
${\mathcal{L}}_{\mathbf{X}}\left(\boldsymbol{\Theta}\right)\xrightarrow{k\to\infty}0\quad\forall\boldsymbol{\Theta}\neq\boldsymbol{\Theta}_{0}.$ (39)
The result in (39) holds for any prior $p(\boldsymbol{\Theta})$ since the prior can be considered a constant with respect to $k$.
Given (39) and the consistency of $\hat{\boldsymbol{\Theta}}$ in Lemma 1, we conclude that [cf. (10) in the main paper]